OpenStudy (anonymous):

Can someone please explain how work done on a gas is equal to -Pi * Vi * ln(Vf / Vi) ?

OpenStudy (anonymous):

This is one of our class slides. I'm confused at how $$ W = -\int P\ dV $$ turns into $$ W = -P_i V_i \int \frac{dV}{V} $$

OpenStudy (anonymous):

Just take a calm look at the question again, you see it gives the gas law pv =nRT ?

OpenStudy (anonymous):

Yes, I see that it gives the gas law. Pressure * Volume is a constant, and is also equal to the number of moles, times the gas constant R, times the temperature.

OpenStudy (anonymous):

I know it's right there, and it's probably extremely obvious.

OpenStudy (anonymous):

Just having trouble seeing it right away.

OpenStudy (anonymous):

Ok, the gas law is true in general for an ideal gas, but your question considers the special case of an Isothermal change, you see that ?

OpenStudy (anonymous):

And so that is why we can say, for these isothermal changes, that PV is constant. - ok so far ?

OpenStudy (anonymous):

Okay, so it's constant because in an isothermal change - the temperature remains constant, and therefore n, R, and T are all constant.

OpenStudy (anonymous):

Right. So now, suppose we know p and v at the start of the change, and at some other point in the process, so we have Pi and Vi, the initial values, and we have P and V some time later. What can we say about PV ?

OpenStudy (anonymous):

Sorry, i meant we know Pi and Vi only - what can we say about PV a bit later ?

OpenStudy (anonymous):

PV is still going to equal the same value, but V will increase from 2m^3 to 3m^3, and therefore P has to drop proportionally

OpenStudy (anonymous):

well if as we said above pv is constant, then we know that pv at the start is the same as pv a bit later, in other words PiVi = PV - we agree on that

OpenStudy (anonymous):

So now, look at your first integral - just take a look at it for a moment to remind yourself what it looks like - ok ?

OpenStudy (anonymous):

Agreed

OpenStudy (anonymous):

Ok, the problem is, we don't know what to do with P - but wait ! what's another expression we can use instead of P ?

OpenStudy (anonymous):

Hint - we were just talking about it.

OpenStudy (anonymous):

I think I see it now! $$ P_i V_i = PV $$ $$ P = \frac{P_iV_i}{V} $$

OpenStudy (anonymous):

yes, i think you got it !

OpenStudy (anonymous):

PiVi is just a constant and can come outside the integral when you replace P with PiVi/V

OpenStudy (anonymous):

then you have to integrate 1/V, which i am sure you know is natural log of V - problem solved ?

OpenStudy (anonymous):

Okay, I understand now that you pointed it out, thank you! But let me ask you - *why* is it that I needed to replace P with PiVi, since P is independent of the integral? Is it because we needed Vi for the integral?

OpenStudy (anonymous):

And yes I have no problems with integration

OpenStudy (anonymous):

Correction: "*why* is it that I needed to replace P with PiVi/V"

OpenStudy (anonymous):

well to do the integral you need to express p in terms of the variable of integration, which in this case is V

OpenStudy (anonymous):

Ahhh, okay, I think I understand now. Thank you for your help!

OpenStudy (anonymous):

welcome